OpenStudy (annie96):

1.)What is the slope of the line that passes through the points (6,4) and (3,8)? 4/3 -3/4 3/4 -4/3 2.)Select the equation of the line that passes through the point (-2.-1) and has slope 5 in point-slope form. (y-1)=5(x-2) (x+2)=5(y-1) (x-2)=5(y+1) (y+1)=5(x+2)

OpenStudy (anonymous):

lets do 2 first point - slope formula looks likr \[y-y_1=m(x-x_1)\] replace \(x_1\) by \(-2\) , \(y_1\) by \(-1\) and \(m\) by \(5\)

OpenStudy (anonymous):

and don't forget that \(y-(-1))=y+1\) etc

OpenStudy (annie96):

(x-2)=5(y+1)

OpenStudy (anonymous):

no

OpenStudy (annie96):

(x+2)=5(y-1)

OpenStudy (anonymous):

\[\huge y-\color{red}{y_1}=\color{blue}{m}(x-\color{green}{x_1})\] \[\huge \color{red}{y_1}=\color{red}{-1}\\ \huge\color{blue}{m}=\color{blue}5\\ \huge \color{green}{x_1}=\color{green}{-2}\]

OpenStudy (anonymous):

Slope of a line fixed by two points: \[points (x_1 , y_1), (x_2 , y_2) \\ y_2 -y_1 =m(x_2-x_1)\]

OpenStudy (annie96):

(y+1)=5(x+2)